2. The accounting department of a large limousine company is analyzing the costs of its services....

2.

The accounting department of a large limousine company is analyzing the costs of its services. The cost data and level of activity for the past 16 months follow.

Month Special Analyses Customer
Accounts
Paychecks
Processed
Accounting
Service Costs
1 4 190 1,150 $ 58,000
2 1 300 1,290 60,800
3 4 240 1,170 58,300
4 4 330 1,150 59,400
5 1 340 970 60,600
6 2 220 1,140 60,500
7 4 200 1,300 59,200
8 3 220 1,020 61,900
9 1 240 900 61,800
10 1 220 1,060 60,300
11 2 270 980 60,700
12 1 210 1,000 59,500
13 2 200 800 60,800
14 3 170 1,260 59,800
15 3 200 850 60,800
16 3 200 1,200 61,900
Totals 39 3,750 17,240 $ 964,300

In addition to the above information, you learn that the accounting department had the following total costs for the past 16 months for each of the following.

Total cost of paychecks processed $ 205,156
Total cost of maintaining customer accounts 113,625
Total cost of performing special analyses 154,050
Total fixed costs (total for 16 months) 491,469
Total costs $ 964,300

Required:

a-1. What is the cost per unit for paychecks processed?

a-2. What is the cost per unit for customer accounts maintained?

a-3. What is the cost per unit for special analyses performed?

b. Assuming the following level of cost-driver volumes for a month, what are the accounting department's estimated costs of doing business using the account analysis approach?

• 1,000 paychecks processed.

• 170 customer accounts maintained.

• 3 special analyses.
Answer #1

A) calculation of per unit costs

1. Paychecks = $205156 / 17240 = $11.9

2. Customer account = $113625/3750 = $30.3

3. Special analysis = $154050/39 = $3950

B) fixed cost per month = $491469/16 = $30716.8

Estimated costs

= $30716.8 + (1000*$11.9) + (170*$30.3) + (3*$3950)

= $59617.8